Our work in locus E502, within unexcavated portions of the 2003 shrine exposure, continues today as does our clearing in E510. The base of locus E504 is producing several LB sherds today as well as the EB materials. At about 470S of the north trench wall the destruction level E514 seems to peter out. Now it's spreading south to envelope the locus 501 area. As we continue to excavate in E514 additional rock platforms / flat stones appear. We also took a soil sample from locus E514 today - of the black, ashy deposits.